### v1

# Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations–B Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations

Abstract
The Twinkle Space Telescope is a satellite designed for spectroscopic observations of a wide range of extrasolar and solar system objects. Equipped with a 0.45 m diameter telescope and a spectrometer covering from 0.5 to 4.5 m simultaneously, Twinkle will be launched in a sun-synchronous, low-Earth orbit, and it is expected to operate for seven years. Twinkle is developed, managed and operated by Blue Skies Space (BSSL), a space science data company whose vision is to accelerate and expand the availability of new, high-quality datasets to researchers worldwide, complementing the space-observatories delivered by government space agencies.
Over its life-time, Twinkle will conduct large-scale survey programs. The scientific objectives and observational strategy of these surveys are defined by researchers who join the Science Team.
Leveraging advances made possible by recent observations with the James Webb Space Telescope, we present here updated simulations evaluating Twinkle’s observational capabilities in the context of exoplanet atmospheres. Through retrieval analyses of HD 209458 b, WASP-107 b, GJ 3470 b, and 55 Cnc e, we demonstrate how increasing observational investment enhances the retrieval of atmospheric parameters and molecular abundances. Our sensitivity study highlights Twinkle’s capability to detect less abundant/detectable molecules depending on the observing strategies adopted. This work provides practical guidance for developing targeted observational strategies to maximise Twinkle’s scientific return.
keywords:
instrumentation: spectrographs – techniques: spectroscopic – planets and satellites: atmospheres – planets and satellites: gaseous planets – planets and satellites: terrestrial planets1 Introduction
The field of exoplanetary science has experienced unprecedented growth, with the number of confirmed exoplanets now exceeding 5,800. A significant proportion of these detections – approximately 4,000 – has been made using transit photometry from space-based surveys such as Kepler, K2, and TESS (Borucki et al., 2010; Howell et al., 2014; Ricker et al., 2014). Since the retirement of the Kepler/K2 mission in 2018, the TESS mission, operational since 2018, has significantly advanced planet discovery efforts. As of 12 March 2025, TESS has identified 7,525 TESS Objects of Interest (TOIs). After excluding false positives, 5,253 candidates remain confirmed by the TESS Follow-up Observing Program Working Group (TFOPWG) (Akeson & Christiansen, 2019). This wealth of discoveries of transiting planets around bright stellar hosts has been a game changer for atmospheric characterisation studies using transit, eclipse and phase-curve spectroscopy.
The European Space Agency’s PLATO, due for launch in 2026, will aim at detecting small transiting planets (above ) around bright stars (11 mag), including terrestrial planets in the habitable zone of solar-like stars (Rauer et al., 2025). Similarly, the Earth 2.0 (ET) mission in transit mode will continuously monitor over 2 million FGKM dwarfs in the original Kepler field and its neighbouring fields for four years to search for new planets, including terrestrial-like planets, across a wide range of orbital periods (Ge et al., 2022, 2024).
Since the first atomic, ionic and molecular species were revealed in exoplanetary atmospheres (e.g., Charbonneau et al., 2002; Vidal-Madjar et al., 2003, 2004; Redfield et al., 2008; Tinetti et al., 2007; Swain et al., 2008, 2009; Grillmair et al., 2008), the interest in studying the chemistry and structure of exoplanet atmospheres has progressively become more widespread. As a result, tens of exoplanetary atmospheres have been observed with the Spitzer and Hubble Space Telescopes and with ground-based facilities in the past couple of decades. These measurements have guided our initial understanding of the atmospheric composition, dynamics and thermal and scattering properties for a variety of atmospheres in very diverse environments (e.g., Knutson et al., 2007; Harrington et al., 2006; Majeau et al., 2012; Stevenson et al., 2014; Demory et al., 2016).
In particular, the WFC3 G141 instrument, covering 1.1 to 1.7 microns, has been central to the hunt for key molecular species such as H2O, CH4, TiO, VO and clouds for a long list of very diverse exoplanets (e.g., Windhorst et al., 2011; Berta et al., 2012; Deming et al., 2013; Fraine et al., 2014; Tsiaras et al., 2016; Arcangeli et al., 2018).
While this effort has been pursued through a number of complementary techniques, including direct imaging and high-resolution spectroscopy from the ground (Bean et al., 2010; Snellen et al., 2010; Nortmann et al., 2018; Macintosh et al., 2015; Chauvin et al., 2017; GRAVITY Collaboration et al., 2019), transit, eclipse, phase-curve spectroscopy and multi-band photometry from space observatories have enabled large-scale atmospheric studies throughout the years (Sing et al., 2016; Tsiaras et al., 2018; Edwards et al., 2023; Changeat et al., 2022; Saba et al., 2025; Dang et al., 2025).
The recent launch of the James Webb Space Telescope (JWST) has offered unprecedented insights into exoplanet atmospheres, spanning wavelengths from 0.5 to 14 m (e.g., JWST Transiting Exoplanet Community Early Release Science Team et al., 2023; Madhusudhan et al., 2023; Xue et al., 2024; Carter et al., 2024; Hu et al., 2024; Bell et al., 2024). JWST has refined measurements of molecular abundances, thermal profiles, atmospheric dynamics, and cloud/haze characteristics. Despite these significant advances, JWST’s broad scientific remit – including galactic and extragalactic astrophysics and solar system studies – limits its dedicated observation time for exoplanetary research.
Addressing this limitation, dedicated missions such as ESA’s Ariel (scheduled for launch in 2029) have been developed. Ariel aims to conduct a comprehensive chemical survey of exoplanetary atmospheres using transit spectroscopy over the spectral range of 0.5–7.8 m, thereby deepening our understanding of how atmospheric chemistry correlates with planetary characteristics (Tinetti et al., 2018, 2022).
The Chinese Space Station Telescope (Xuntian), scheduled for late 2026, will offer high-resolution multiband imaging and slitless spectroscopy across 0.255–1 m (Zhan, 2021).
The Twinkle Space Telescope, managed and operated by Blue Skies Space Ltd. (BSSL), will be dedicated to provide spectroscopic data for various science themes developed by its members (Jason et al., 2016; Savini et al., 2018; Edwards et al., 2019a, b; Archer et al., 2020; Stotesbury et al., 2022, 2024). Equipped with a 0.45-metre telescope and a spectrometer covering simultaneously from 0.5 to 4.5 m, Twinkle is capable of providing spectra for a wide range of targets such as brown dwarfs, stars, protoplanetary disks, Solar System objects and exoplanets. Two large survey programmes focusing on extrasolar targets and Solar System objects are being planned during the first three years of Twinkle’s expected seven years operations (Stotesbury et al., 2022). While BSSL is responsible for the satellite, the data analysis tools and facilitating the survey planning process, researchers who joins the survey program (Science Team) will decide the scientific objectives and observational strategy. Twinkle Science Team members have been tasked with delivering a comprehensive and scientifically optimised observation plan and list of targets.
Based on the existing scientific themes proposed by the Twinkle Science Team members, a large fraction of the survey time will be allocated to extrasolar objects and the atmospheric characterisation of diverse types of exoplanets is expected to be a significant part of the survey. This paper presents new simulations to evaluate Twinkle’s potential for atmospheric characterisation in the context of more recent findings by JWST and discusses possible approaches to optimise the observational strategy.
Many of these upcoming missions have spectral ranges that significantly overlap with JWST, making JWST’s observations an invaluable reference for simulating their expected performance. The combination of JWST’s comprehensive spectral coverage and its recent discoveries offers essential benchmarks that drive our updated simulations of Twinkle’s observational capabilities. In this study, we incorporate the latest JWST-calibrated atmospheric models into Twinkle’s simulation framework, leading to more accurate predictions of signal-to-noise ratios (SNR) and a refined assessment of the detectability of key atmospheric features.
Previous studies by Edwards et al. (2019c) evaluated Twinkle’s expected performance to study the atmospheric composition and structure of three iconic exoplanets, HD 209458 b, GJ 3470 b, and 55 Cnc e. However, due to lack of IR spectroscopic data pre-JWST, those analyses were based on theoretical models and simplified assumptions. Recent JWST observations have revealed complexities in atmospheric chemistry, thermal and cloud structures that necessitate an update to these models (e.g., Xue et al., 2024; Beatty et al., 2024; Hu et al., 2024). More specifically, recent JWST observations of HD 209458 b by Xue et al. (2024) have provided a high-quality transmission spectrum from 2.3 to 5.1 m, revealing pronounced absorption features from CO2 and H2O. Based on thermochemical equilibrium assumptions, these results suggest a supersolar metallicity and a low C/O ratio (0.11). GJ 3470 b has emerged as a key target of interest due to its relatively transparent atmosphere and the detection of SO2 in its atmosphere, reinforcing the evidence of disequilibrium chemistry in lower-mass exoplanet atmospheres (Beatty et al., 2024). JWST’s eclipse observations of the super-Earth 55 Cnc e by Hu et al. (2024) have confirmed the presence of a volatile-rich atmosphere, potentially maintained by a magma ocean. In addition to these three exoplanets, we have added here the study of WASP-107 b: Welbanks et al. (2024) have identified the presence of several molecular species in the atmosphere of this planet, including H2O, CH4, CO, CO2, SO2, and NH3 when combining HST and JWST data.
| HD 209458 b | WASP-107 b | GJ 3470 b | 55 Cancri e | |||||||||
| Planet Type | Hot Jupiter | Warm Neptune | Warm Neptune | Super Earth | ||||||||
| Planet Mass () | ||||||||||||
| Planet Radius () | ||||||||||||
| Teq (K) | ||||||||||||
| Stellar Magnitude (V/K) |
| JWST Availability (m) | 2.3–4.5 | 2.4–4.5 | 2.45–4.97 | 3.94–12 | ||||||||
| HST Wavelength Range (m) | 0.5–0.9; 1.1–1.7 | 0.8–1.6 |
This paper first reviews Twinkle’s instrumental capability, the models and methodology used in estimating Twinkle’s science performance in Section 2. It updates the radiometric model employed to simulate the noise characteristics (Section 2.1) and the list of potential target candidates within Twinkle’s field of view and details. Section 2.3 describes the atmospheric forward and inverse models obtained with the open-source exoplanet atmospheric retrieval framework TauREx 3. Simulations and retrieval results are presented in Section 3, highlighting Twinkle’s ability to detect molecular species and other atmospheric properties across a diverse exoplanetary sample and observational strategies, based on recent JWST advances. Finally, Section 4 discusses the results, addresses study limitations, and proposes directions for future simulation work.
2 Methods
2.1 Twinkle’s Radiometric Model
Twinkle will observe spectra simultaneously across two channels covering the wavelength range from 0.5 to 4.5 m. Channel 0 operates from 0.5 to 2.43 m with a resolving power () of up to 70, and channel 1 operates from 2.43 to 4.5 m with a resolving power of up to 50.
Twinkle’s radiometric performance is assessed using the Twinkle Radiometric Tool (Stotesbury et al., 2022), adapted from the open-source radiometric software ExoRad (Mugnai et al., 2023). The radiometric model simulates observational and instrumental noise contributions, including photon noise, instrument emission, detector dark current, read noise, and zodiacal background. This model has been refined to incorporate recent advances in instrument design. It calculates realistic noise estimates, which in turn allows determination of the number of observations necessary to reach a specific SNR through observation stacking. The cumulative SNR achieved by combining multiple observations can be expressed as:
where is the number of observations, and is the signal-to-noise ratio (SNR) of a single observation. The signal is derived from the amplitude () of spectral features in transit spectra, approximated by:
Here, and represent the planetary and stellar radii, respectively, and denotes atmospheric thickness. The atmospheric thickness is typically estimated as , with ranging from 3 to 5 for hydrogen/helium (H2/He)-dominated atmospheres. The scale height is defined by:
where is Boltzmann’s constant, the equilibrium temperature of the planet, the planet’s surface gravity, and the mean molecular weight of the atmosphere (usually assumed as 2.3 for H2/He atmospheres but refined using recent JWST data in this study).
Previously, due to limited observational constraints, atmospheric thickness () was set as 1, 3, or 5 scale heights, with 5 scale heights being a commonly used assumption (e.g., Edwards et al., 2019c; Booth et al., 2024). However, with the aid of recent data release from JWST and its overlapping wavelength coverage with Twinkle, constraints on atmospheric composition have improved, enabling more precise estimates of atmospheric scale height and thickness above the optically thick continuum. Thus, we can derive a more realistic SNR and resolution needed for Twinkle to better retrieve the atmospheric chemical composition, and consequently, determine the number of observations required to achieve a given SNR. It is worth noting that the radiometric model employed in this study calculates Twinkle’s observational errors assuming complete transit/eclipse coverage. However, observational availability of targets are influenced by Earth obstruction due to Twinkle’s orbital configuration, causing Twinkle to sometimes capture partial transits or eclipse events, similar to the situation encountered by CHEOPS (Benz et al., 2021).
While for some exoplanets, the number of required observations is high, it is important to note that lowering spectral resolution can increase the SNR per wavelength bin and thus reduce the observational requirements. Although this method is widely used (e.g., Kreidberg et al., 2014; Stevenson et al., 2014), artificially reducing resolution in simulations introduces bias along the wavelength axis, particularly around spectral peaks. To prevent this bias, this study retrieves spectra at Twinkle’s full resolution. After the launch of Twinkle, the actual resolution adjustment will effectively improve this bias.
2.2 Twinkle’s Target Candidates
Twinkle will be placed in a 1200 km, sun-synchronous low-Earth orbit, with a 6am Local Time of Ascending Node (LTAN). The spacecraft’s observational capabilities are restricted by the Field of Regard (FoR) (a 40 degree cone around the anti-Sun vector), solar & lunar exclusion constraints and Earth obstruction across different seasons.(Stotesbury et al., 2022).
To systematically assess target observability, we employed the BSSL Twinkle Radiometric Tool to simulate observational noise for all confirmed exoplanets within Twinkle’s FOR possessing well-defined stellar and planetary parameters. Planetary and stellar data from the NASA Exoplanet Archive are taken as input into the radiometric tool to simulate noise levels for single transits or eclipses at Twinkle’s maximum resolving power. For TESS Objects of Interest (TOIs) lacking direct planetary mass measurements and thus not able to be processed with Radiometric Tool, masses were estimated using the methodology described by Chen & Kipping (2017). Approximately 200 TOIs within Twinkle’s FOR were subsequently excluded due to unavailable stellar mass data. Additionally, initial mass-radius relationships for gas giants between 8 and 22 Earth radii proved inconsistent with current empirical data. Hence, planetary masses for these TOIs were recalculated based on a log-normal distribution fitted to known transiting exoplanet populations, as detailed in Figure 15 and further described in Appendix B.
Our primary focus was on exoplanets with H2/He dominated atmospheres; thus, we assumed a mean molecular weight of 2.3 for initial assessments. Atmospheric thicknesses equivalent to 3 atmospheric scale heights were adopted uniformly for all candidate targets. Exoplanets predicted to achieve median SNR levels of 3, 5, or 7 within 30 transits or eclipses at Twinkle’s native spectral resolution were considered suitable candidates for detailed atmospheric characterisation.
2.3 Simulated Exoplanets and Spectral Retrievals
The detection of molecular species and characterisation of chemical compositions and abundances in exoplanet atmospheres remain central objectives of transit spectroscopy. Achieving these goals offers critical insights into planetary processes, formation histories, and potential habitability. Twinkle’s extensive and continuous spectral coverage from optical to near-infrared wavelengths includes absorption features for many previously detected and theoretically predicted molecules. In this study, we conduct atmospheric retrieval simulations on selected exoplanets (listed in Table 1), chosen from Twinkle’s candidate list based on the availability of JWST data. These planets represent a diverse range of types, aligning with Twinkle’s broader exoplanetary science themes, and facilitating a comprehensive evaluation of Twinkle’s capabilities informed by recent JWST findings.
We employed the open-source retrieval code TauREx 3.2.2 (Al-Refaie et al., 2021) to simulate forward models for transmission and emission spectra and to perform spectral retrieval analyses. Stellar spectra for each target star were simulated using the PHOENIX stellar library (Husser et al., 2013). Chemical profiles were generated using the FastChem equilibrium chemistry plugin for equilibrium assumptions, supplemented by TauREx’s free chemistry feature for non-equilibrium chemistry scenarios. All atmospheric models assumed a plane-parallel geometry comprising 100 layers, including molecular absorption, Rayleigh scattering, grey clouds, and collision-induced absorption (CIA). For all exoplanets, CIA processes involving H2-H2 and H2-He were considered, with additional contributions from N2-N2 and O2-O2 specifically included for 55 Cnc e. Molecular opacity cross-sections were sourced from the ExoMol database (Tennyson et al., 2024), while CIA data were taken from the HITRAN database (Karman et al., 2019).
Retrieval analyses were performed using the Bayesian MultiNest algorithm (Feroz et al., 2009), employing 1000 live points and an evidence tolerance threshold of 0.5. Non-informative priors for molecular abundances were uniformly set on a log scale within the range of to , while other planetary parameters utilised priors constrained by previous literature of specific planets.
To incorporate recent JWST observational findings, chemical compositions and planetary parameters for each simulated exoplanet were based on best-fit solutions reported in recent literature, alongside equilibrium chemistry predictions. Forward modelling and retrieval analyses were performed based on simulated transmission and emission spectra incorporating predicted Twinkle observational errors.
For HD 209458 b, recent JWST observations by Xue et al. (2024) were combined with HST data from Saba et al. (2025). The simulated spectrum is fitted with observed data through minor adjustments. Similarly, to assess Twinkle’s capabilities for WASP-107 b, a simulated spectrum covering Twinkle’s wavelength coverage (0.5–4.5 m) was constructed using HST and JWST NIRSpec data as benchmarks. The atmospheric model included known molecular constituents identified by Sing et al. (2024), specifically H2O, CH4, NH3, CO, CO2, SO2, and H2S, alongside potential SiO detection proposed by Ma et al. (2025), and optical absorbers Na and K, with uniform abundances assumed.
For GJ 3470 b, we simulated Twinkle’s performance by closely aligning spectral data with JWST and HST observations across the 0.5–4.5 m range. The atmospheric composition model incorporated nine primary molecules (H2O, CH4, CO2, CO, SiO, SO2, NH3, HCN, H2S), with atmospheric temperature fixed at 600 K and a deep cloud deck placed at 3000 Pa to highlight molecular absorption signatures. For the emission spectra of 55 Cnc e, we employed atmospheric assumptions consistent with Hu et al. (2024), featuring predominant CO2 and CO composition alongside trace species (C2H2, HCN, O2, and N2) with constant mixing ratios. Given JWST’s spectral coverage starting around 4 m, only select JWST NIRCam data points were used as references for the simulated Twinkle spectra.
To evaluate the retrievability of minor atmospheric species whose spectral signatures are typically obscured by clouds or dominated by major molecules, we conducted a dedicated sensitivity study. Using WASP-107 b as the test case, we systematically amplified the abundances of minor species (CH4, SO2, NH3, and H2S) to identify the enhancement factor necessary for their spectral features to become clearly observable above simulated observational uncertainties. The baseline scenario employed the previously established retrieval setup for WASP-107 b, stacking six transits to achieve a median SNR of 10. Individual retrieval analyses were then performed separately for each molecule, progressively increasing their abundances, to quantify their retrieval thresholds under varying observational conditions.
3 Results
3.1 Model assumptions: lessons learned from JWST
With recent advances resulting from JWST observations significantly refining our understanding of exoplanetary atmospheres, previous assumptions regarding atmospheric thickness and composition require reassessment. JWST data have provided critical updates, particularly regarding atmospheric composition and the prevalence of clouds, which directly impact observational strategies and the SNR achievable by Twinkle. Table 2 summarises our revised estimates of the number of transits required by Twinkle to reach specific SNR targets (3, 5, 7, and 10) at native resolution after incorporating recent JWST findings. We compare these updated estimates with earlier calculations that assumed atmospheric thicknesses of 2, 3, and 5 scale heights. The updated calculations assume ideal observational conditions – complete transit visibility without interruptions – thus serving as lower-bound estimates. These preliminary outcomes, integrating JWST observational results, will continue to evolve as additional JWST data become available.
| Planet | Median SNR | Predicted Transits (5) | Predicted Transits (3) | Predicted Transits (2) | Realistic Transits |
| HD 209458 b | 5 | 1 (0.18) | 1 (0.51) | 2 (1.21) | 1 (0.85) |
| 7 | 1 (0.35) | 2 (1.04) | 3 (2.6) | 2 (1.77) | |
| 10 | 1 (0.75) | 3 (2.32) | 7 (6.79) | 5 (4.24) | |
| WASP-39 b | 5 | 3 (2.26) | 7 (6.31) | 15 (14.31) | 3 (2.11) |
| 7 | 5 (4.45) | 13 (12.45) | 29 (28.45) | 5 (4.14) | |
| 10 | 10 (9.12) | 26 (25.74) | 60 (59.84) | 9 (8.48) | |
| WASP-80 b | 5 | 3 (2.01) | 6 (5.7) | 14 (13.41) | 8 (7.49) |
| 7 | 4 (3.98) | 12 (11.56) | 29 (28.5) | 16 (15.35) | |
| 10 | 9 (8.33) | 26 (25.44) | 71 (70.86) | 35 (34.67) | |
| WASP-107 b | 5 | 1 (0.16) | 1 (0.45) | 2 (1.02) | 2 (1.31) |
| 7 | 1 (0.32) | 1 (0.88) | 2 (2.00) | 3 (2.59) | |
| 10 | 1 (0.65) | 2 (1.81) | 5 (4.15) | 6 (5.4) | |
| GJ 3470 b | 3 | 1 (0.59) | 2 (1.65) | 4 (3.77) | 4 (3.58) |
| 5 | 2 (1.64) | 5 (4.69) | 12 (11.09) | 11 (10.49) | |
| 7 | 4 (3.27) | 10 (9.55) | 24 (23.85) | 23 (22.45) |
The realistic observation requirements based on recent JWST constraints are marked in red in Table 2, while atmospheric thickness predictions closest to the JWST-derived reality are highlighted in green. Initial analyses indicate that Jupiter-sized exoplanets (e.g., HD 209458 b, WASP-39 b, WASP-80 b) typically have observable atmospheric thicknesses ranging from approximately 2.5 to 5 scale heights. In contrast, Neptune-sized exoplanets (e.g., WASP-107 b, GJ 3470 b) align better with atmospheric thicknesses around two scale heights. These findings broadly confirm previous assumptions but now offer significantly tighter constraints, improving observational efficiency.
Furthermore, JWST data confirm that cloud coverage significantly reduces observable atmospheric thickness. Most H2/He-rich planets studied thus far display observable atmospheric features between 2 and 3 times their scale height. Notably, exceptions such as WASP-39 b exhibit significantly extended atmospheric signals, a result directly linked to its particularly low planetary density. These refined assumptions provide a robust foundation for the detailed retrieval studies and sensitivity analyses of individual exoplanets presented in subsequent sections.
3.2 Candidate List for Atmospheric Characterisation Survey
Among the confirmed exoplanets, approximately 1,070 transiting exoplanets lie within Twinkle’s FOR. Of the TOIs, 2,009 candidates fall within Twinkle’s FOR. Figure 1 illustrates the right ascension (RA) and declination (Dec) of all confirmed exoplanets and TOIs distribution in the Twinkle’s FOR.
After processing all targets through the radiometric model, a combined set of known exoplanets and TOIs capable of achieving median SNR thresholds (3, 5, or 7) within 30 observations at native resolution is presented in Figure 2.
This initial candidate list used for the observability study for Twinkle’s atmospheric characterisation survey is specifically tailored to H2/He-rich planets, due to current uncertainties surrounding the atmospheric composition of smaller planets. Recent observations of terrestrial planets have highlighted significant diversity, emphasising the need for further study with dedicated future missions. Consequently, terrestrial planets have not been extensively analysed in this observability study. Further refinement of the candidate list will be informed by complementary observations from JWST, PLATO, CSST, and Earth 2.0, enabling tailored observational strategies aligned with the scientific objectives of the Twinkle Science Team.
3.3 Spectral Retrievals: Gas Giants
3.3.1 HD 209458 b
Simulated spectra for HD 209458 b with Twinkle’s predicted uncertainties after fitting with combined set of HST WFC3 and JWST are shown in Figure 3. Retrievals conducted under chemical equilibrium assumptions indicate that, even with a single transit, abundances of major molecules can be accurately recovered. However, minor molecular species such as CH4, SO2, and HCN exhibit weaker constraints, introducing uncertainties in metallicity and C/O ratio estimates (Figure 4). Increasing observations to 10 transits significantly enhances the accuracy of CH4 and HCN abundance constraints, leading to more robust metallicity and C/O ratio estimates (Figure 4). This improvement mainly results from the distinct and abundant H2O absorption due to the assumed low C/O ratio, prominent TiO features at optical wavelengths, and clearly defined CO2 features at the limit of Twinkle’s coverage. Additional planetary parameters, including radius, atmospheric temperature, and cloud-top pressure, were also accurately recovered within .
To provide a more comprehensive assessment, we performed a free retrieval analysis without assuming chemical equilibrium by selecting the eleven most critical molecules identified in previous simulations, along with species suggested by Xue et al. (2024). A constant vertical mixing ratio (VMR) profile based on equilibrium abundances and suggestions from Xue et al. (2024) was used for these molecules in a free retrieval analysis. Results in Figure 11 confirm robust retrieval of H2O, TiO, and VO, whereas other molecular signatures remain masked by the dominant H2O absorption and cloud opacity. Twinkle’s extensive spectral coverage from optical to infrared allows for clear identification of TiO and VO absorption features. However, due to the relatively low assumed abundance of CO2 () and increased noise at longer wavelengths, the CO2 feature was only partially detected and thus less reliably constrained.
Retrieval comparisons at different SNR values (5, 7, and 10) for HD 209458 b, as illustrated in Figure 11, indicate improved parameter constraints for strongly featured molecules such as H2O and TiO at higher SNR. Conversely, parameters not adequately retrieved at an SNR of 5 exhibit limited improvement at higher SNR levels. This limitation of improving SNR and difficulty of retrieving minor species in HD 209458 b likely arises from spectral masking by prominent water absorption and simplified grey cloud assumptions, rather than reflecting Twinkle’s intrinsic capabilities.
| Parameter | Truth | Retrieved (SNR = 5) |
| Planet Radius [] | 1.35 | |
| Temperature [K] | 1140 | |
| Detection | ||
| Upper Limit | ||
| 2.11 | ||
3.3.2 WASP-107 b
Simulated spectrum for WASP-107 b with Twinkle’s predicted error after fitting with available HST WFC3 and JWST NIRCam from Welbanks et al. (2024) are shown in Figure 5. Calculations shown in Table 2 suggest the atmospheric thickness of WASP-107 b is slightly less than two scale heights, likely due to high-altitude cloud cover. To reflect this observational constraint, our simulation adopted a high cloud altitude at 70 Pa.
Posterior distributions for WASP-107 b in Figure 12 revealed significant degeneracies between the planetary radius and cloud-top height parameters, with the radius typically underestimated and cloud height overestimated in retrievals. Nevertheless, robust constraints were obtained for the equilibrium temperature and the abundances of H2O, CO2, and Na. The abundances of NH3 and CH4 remainundetectable due to their lower contribution compared with cloud opacity within Twinkle’s spectral coverage as in Figure 5. The primary absorption band for CO (4.5–5.0 m) lies outside Twinkle’s spectral range, making CO undetectable. The spectral signature of SO2 at 4.0 m is too weak to be reliably retrieved, primarily due to interference from cloud opacity and observational noise.
Comparative retrieval results at different SNR values showed that parameters clearly observable in the spectra, such as H2O and CO2, exhibited tighter constraints with increasing SNR. Parameters initially unconstrained at an SNR of 5 but exhibiting spectral contributions stronger than cloud opacity—such as Na—became retrievable at higher SNR values (SNR = 10). Furthermore, while CO2 could be constrained at all simulated SNR levels, increasing the SNR to 7 or 10 significantly improved the accuracy and reliability of its retrieval.
| Parameter | Truth | RP (SNR = 5) |
| Planet Radius [] | 0.933 | |
| [K] | 700 | |
| Detection | ||
| Upper Limit | ||
3.3.3 GJ 3470 b
Simulated spectra of GJ 3470 b with Twinkle’s predicted error after fitting with available HST WFC3 and JWST NIRCam from Saba et al. (2025) and Beatty et al. (2024) are shown in Figure 6. The contribution plot illustrates distinct spectral features from key molecules. In addition to strong absorption by H2O, significant CO2 features at 2.0, 3.8, and 4.4 m allow robust constraints on its abundance. Prominent CH4 absorption features at 1.6, 2.4, and 3.4 m are also clearly discernible. A detectable but weaker absorption peak of SO2 at 4.0 m provides crucial insight into atmospheric disequilibrium chemistry.
Considering the observational challenges posed by GJ 3470 b’s low transit depth and relatively flat spectrum, we evaluated retrieval outcomes at SNR levels of 3, 5, and 7, approximately 23 transits are required to reach an SNR of 7. The retrieval results, depicted in Figure 13, demonstrate accurate constraints for parameters such as planetary radius and equilibrium temperature. Conversely, the cloud deck height remains unconstrained due to its depth, rendering it indistinguishable from other atmospheric components. The abundances of H2O, CH4, and CO2 were successfully retrieved at all SNRs considered. However, SO2 could only be effectively constrained at the highest SNR (SNR = 7). Minor molecular species such as CO, SiO, NH3, HCN, and H2S were undetectable even at high SNR due to their weak absorption features within Twinkle’s spectral range.
Comparative retrieval results across different SNR scenarios yielded consistent conclusions with previous exoplanet cases. Parameters with strong spectral signatures above cloud opacity, such as H2O, CO2, and CH4, benefit significantly from higher SNR, resulting in tighter constraints. Conversely, molecules like SO2, with a single prominent absorption feature near the spectral range boundary, require high SNR levels for accurate abundance retrieval.
| Parameter | Truth | RP (SNR = 3) |
| Planet Radius [] | 0.36 | |
| Temperature [K] | 600 | |
| Detection | ||
| Upper Limit | ||
3.4 Spectral Retrievals: super-Earths
3.4.1 55 Cancri e
Figure 7 shows a simulated spectrum for 55 Cnc e, along with Twinkle’s predicted error after fitting with available JWST NIRCam data from Hu et al. (2024). This baseline scenario uses three blackbody curves representing temperatures of 1400, 2000, and 2500 K. Comparing these baselines with the simulated and observed data points reveals distinctive absorption features present in the emission spectrum. Twinkle’s expected performance, calculated by reducing the resolution of channel 0 to one-third and channel 1 to half their native resolutions and stacking ten eclipse observations, demonstrates comparable sensitivity to JWST within its spectral range. A combined analysis of Figures 7 and 8 reveals significant spectral divergence from blackbody emission beginning around 2.0 m. Although the higher resolution of channel 0 below this wavelength is underutilised, the spectral range above 2.0 m effectively captures prominent CO2 absorption features. Additional CO2 features are clearly identifiable within Twinkle’s channel 1 coverage. Figure 8 further highlights that CO2 dominates the spectral signature throughout Twinkle’s wavelength range, overshadowing contributions from minor species such as C2H2, CO, and HCN. Notably, an increase in CO contribution is evident beyond 4.2 m, suggesting enhanced detectability if Twinkle’s coverage could extend slightly further.
Using this simulated spectrum as input, we performed emission retrievals to evaluate Twinkle’s capability in constraining atmospheric properties. To prevent spectral bias from wavelength binning, retrievals were conducted at Twinkle’s native resolution, despite larger resultant uncertainties. Retrieval outcomes for scenarios involving 10, 20, and 30 stacked eclipse observations are presented in Figure 14. The results illustrate that, with just 10 eclipses, Twinkle effectively constrains the temperature structure and reliably retrieves the abundance of the dominant atmospheric species, CO2. Additional observations enhance parameter constraints incrementally but do not significantly improve the detection of dominant atmospheric constituents. This demonstrates Twinkle’s robust performance, particularly for bright terrestrial targets.
| Parameter | Truth | RP (10 eclipses) |
| Planet Radius [] | 0.167 | |
| [K] | 2300 | |
| [K] | 1500 | |
| [Pa]) | 5 | |
| [Pa]) | 2 | |
| Detection | ||
| Upper Limit | ||
3.5 Minor Species Sensitivity Study on WASP-107 b
Contribution plots from simulations (Figures 5, 6, and 8) revealed that several minor species, such as CH4, SO2, NH3, and H2S, typically possess abundances too low to yield distinguishable spectral features, especially in atmospheres dominated by water vapour absorption or obscured by clouds. Direct retrieval of these species from featureless continuum regions was thus found unreliable, as illustrated in posterior plots provided in the Appendix. Such limitations reflect realistic observational challenges faced when attempting to detect minor constituents even at higher SNR levels.
To better quantify these observational constraints, we assessed the amplification factors necessary for each minor species’ spectral features to surpass the noise threshold of Twinkle observations at SNR = 10 (Comparison of the original and amplified sepctra in Figure 9). Our results show that CH4 and SO2 features become clearly observable when their abundances increase by approximately fivefold, corresponding to abundances of and , respectively. For NH3 and H2S, which have broader and weaker absorption bands overlapping significantly with water vapour, spectral signatures required approximately a twentyfold abundance increase (to and ) before becoming discernible.
Individual retrieval analyses for each amplified abundance scenario confirmed that minor species, initially undetectable at baseline abundances, can indeed be reliably retrieved when their spectral signatures exceed noise thresholds (posterior distributions in Figure 10). At Twinkle’s native resolution with SNR = 10, CH4 and SO2 could be constrained down to abundances of and , respectively. Lowering the SNR to 5 increased these retrieval thresholds to approximately for CH4 and for SO2. NH3 could be reliably constrained at an abundance of at SNR = 10, requiring further amplification to (fortyfold increase from baseline) at SNR = 5. H2S retrieval was challenging at SNR = 5 due to its weaker, indistinct absorption features compared to water.
The goal of this sensitivity analysis was to isolate the impact of instrumental capabilities (resolution, SNR) from atmospheric factors such as inherently low abundances and optically thick cloud cover. By establishing clear thresholds for the detectability and retrievability of minor atmospheric constituents at varying levels of observational sensitivity, our results provide valuable guidelines for mission planners.
4 Discussion
JWST’s observations provide a critical empirical foundation for simulating expected outcomes with future telescopes, particularly for targets already characterised by JWST. For targets yet to be observed, empirical relationships derived from JWST data – such as those relating to atmospheric mean molecular weight or number of scale heights (see Table 2) – offer valuable guidance for initial observational planning. These preliminary simulations are essential, as they enable mission teams to refine their observational strategies and allocate resources efficiently to best achieve their scientific goals. In this context, our study evaluating Twinkle’s capabilities for atmospheric characterisation offers important initial insights into the mission’s anticipated performance in exoplanet surveys, leveraging the most recent discoveries from JWST. Our results demonstrate that, through the use of low-resolution spectroscopy and strategies involving the stacking of multiple observations, Twinkle’s observations will effectively allow to determine planetary parameters and key chemical components for hundreds of exoplanets.
Focusing on transit spectroscopy of H2/He-dominated atmospheres, we found that when a median SNR of 3 is achieved, key planetary parameters such as radius, atmospheric temperature and cloud heights can be constrained within meaningful ranges (e.g., GJ 3470 b shown in Figure 13). For chemical species identifiable in the spectra, retrievability improves significantly as SNR increases. However, higher SNR does not always enhance the detectability of spectral features from molecules that are intrinsically weak or masked by clouds or other more dominant chemical species, emphasising the intrinsic limitations imposed by planetary atmospheres themselves or technique adopted or Twinkle’s spectral coverage and/or resolving power. In eclipse spectroscopy, Twinkle also exhibits promising capabilities, as demonstrated in our simulations for the super-Earth 55 Cnc e (more details in Appendix 3.4.1). While characterising detailed chemical compositions remains challenging, our simulations confirm Twinkle’s ability to extract fundamental planetary parameters, such as the atmospheric thermal structure, critical for investigating internal/surface properties and heat transport processes. Moreover, with thousands of hours of observing time available to Twinkle’s science team, continuous phase-curve observations are feasible, particularly for short-period exoplanets such as 55 Cnc e. Such phase-curve data will offer insights into the three-dimensional structure of planetary atmospheres, complementing other observational approaches (Knutson et al., 2007; Stevenson et al., 2014; Demory et al., 2016; Dang et al., 2025). Additionally, Twinkle’s ability to observe multiple transits and its high-precision timing capabilities will enhance transit ephemeris accuracy and establish more robust baselines for transit timing variation (TTV) analyses. Twinkle’s participation in TTV and transit duration variation (TDV) studies will further refine orbital parameters, improve planet mass determinations in multi-planet systems, and aid in detecting potential new exoplanets.
4.1 H2/He dominated exoplanets
Simulations presented for gas giants and Neptunes have significantly benefited from recent JWST observations, allowing updated assessments of their observability with Twinkle through improved understanding of their atmospheric compositions. For bright hot Jupiters, such as HD 209458 b, Twinkle can achieve an SNR of 10 by stacking just 5 transits at native resolution. As illustrated in Figure 11, the improvement in retrievability for planets with prominent spectral features becomes less significant at higher SNRs. This diminished improvement for HD 209458 b arises because the error associated with a single transit observation by Twinkle is already relatively small. Consequently, while further enhancements through additional observations yield limited benefits, bright targets offer an excellent opportunity to study stellar and/or atmospheric variability. For bright H2/He-dominant exoplanets without prior data, an efficient strategy would be to initially perform a single transit and/or eclipse observation, conduct preliminary analysis, and then decide whether further observations are required based on initial findings or existing data from other instruments.
In contrast, for warm H2/He-dominant exoplanets orbiting fainter stars, such as WASP-107 b, lower intrinsic observability leads to larger errors in simulated observations. Thus, increased stacking of transits and corresponding SNR improvements markedly enhance retrievability. Retrieval results for WASP-107 b demonstrate that weaker species like Na become detectable only at higher SNR (e.g., SNR = 10). For this type of targets, an initial estimate of the required number of transits based on prior knowledge and scientific objectives would be advisable. If there is no previous knowledge, a good approach could be to start with a limited number of transits and then performing an initial analysis on the data acquired to design the follow-up observational plan.
Neptunes and sub-Neptune-sized planets are often cloudy and/or their atmospheres are heavier than hydrogen. For these reasons, typically ten to twenty transits are required for adequate SNR, the lower intrinsic signal results in greater gains when improving from lower to moderate SNR. For instance, GJ 3470 b parameters were weakly constrained at SNR = 3 but became well constrained at SNR = 5 for molecules with multiple distinct spectral features. However, retrieval of species such as SO2, having limited spectral signatures, required higher SNR (SNR = 7). Increasing SNR beyond this point yielded diminishing returns. Similar observations were made for WASP-107 b, where increasing observations from SNR = 5 (2 transits) to SNR = 7 (3 transits) provided more improvement than further increasing to SNR = 10 (6 transits). The diminishing marginal returns, as described by Equation 1, suggest an optimal observational strategy for smaller-radius planets targeting initially median SNR levels (3 or 5). Comparing retrieval improvements at these levels and balancing against the observational investment may help optimise follow-up observations.
In summary, Twinkle’s capability to characterise exoplanet atmospheres depends primarily on stellar brightness in optical and infrared wavelengths, planetary and stellar radii, main atmospheric component and cloud coverage. While stellar brightness and planetary dimensions are usually well-characterised by previous observations, cloud properties and, in the case of small planets, atmospheric mean molecular weight may require dedicated characterisation efforts. Cloud contributions may significantly impact parameters’ retrieval, especially for minor species. For instance, in WASP-107 b, SO2 spectral features were fully obscured by cloud opacity, and there are degeneracies between cloud height and planetary radius. Incorporating detailed cloud microphysical and radiative models into future simulations might allow more realistic analyses.
4.2 Planets with Non H2/He Dominated Atmosphere
Atmospheric characterisation of super-Earths presents significant challenges due to their relatively small radii and heavier atmospheres, resulting in weaker observable signals. Eclipse simulations for 55 Cnc e suggest that the primary information expected from such planets is their temperature structure and constraints in atmospheric composition, distinguishable through comparisons with blackbody emission curves. If an atmosphere exists and is dominated by detectable molecules such as CO2, our simulations indicate that Twinkle can well constrain those. For bright targets such as 55 Cnc e, increasing the number of eclipses beyond an initial set (10 eclipses for 55 Cnc e) shows diminishing returns in parameters’ retrievability. An alternative approach for gaining deeper insights into super-Earths is phase-curve spectroscopy. The short orbital period of 55 Cnc e makes continuous phase-curve monitoring feasible. We are currently conducting simulations to assess Twinkle’s potential in phase-curve observations, and these findings will be detailed in future studies.
4.3 Stellar Activity Effects, Targets Observability and Future Works
While stacking multiple observations to achieve higher SNR per wavelength bin is an effective strategy to enhance target observability, variability between observations introduces potential challenges. Variations in planetary atmospheric conditions between different transits or eclipses can lead to differences in the collected spectra. More significantly, stellar activity may significantly contaminate atmospheric parameters such as molecular abundances and temperature retrievals. A population study of 20 exoplanets conducted with HST by Saba et al. (2025) demonstrated that stellar contamination could alter molecular abundance estimates by up to six orders of magnitude and atmospheric temperatures by up to 145%. Stellar contamination predominantly affects the optical wavelengths, which are partially covered by Twinkle. Hence, the potential influence of stellar activity must be carefully considered, especially when multiple observations are stacked. Investigating stellar activity effects on exoplanet atmospheres is one of Twinkle’s core scientific themes, and a future study will explore Twinkle’s capability to identify and quantify these stellar variations.
Another significant constraint is observational availability, influenced by Earth obstruction due to Twinkle’s orbital configuration, planetary orbital periods, and transit durations. To address this, the Twinkle team has developed an orbital planning tool, allowing members to identify available observational windows over a year. Integrating this availability with Twinkle’s simulated performance enables the development of optimised, well-informed observation strategies for target exoplanets.
However, one remaining problem is that Earth obstruction also causing Twinkle to sometimes capture partial transit or eclipse events, as described in Section 2.1. Consequently, a more realistic estimate of the number of required observations should incorporate an observing efficiency factor, as demonstrated by Booth et al. (2024). Meanwhile, the Twinkle team is actively developing methodologies for effectively stacking multiple partial transits or eclipses to maximise SNR. Comparable techniques have already been demonstrated for combining multiple nights of data with the same instrumental setup in both transmission and emission studies from ground-based instruments (e.g., Giacobbe et al., 2021; Kesseli et al., 2022; Scandariato et al., 2023). A simulation by Boldt-Christmas et al. (2024) specifically addresses the compromises required when maximising SNR through stacked observations for transiting exoplanets.
Several studies underline the importance of combining multiple observations for the future characterisation of Earth-like and super-Earth exoplanets, even for powerful observatories such as JWST (Wunderlich et al., 2019). Indeed, recent analysis by Lustig-Yaeger et al. (2023) involved combining two JWST transit observations for improved results. Historically, enhancing SNR through stacking multiple observations has often been impractical due to limited telescope availability. However, as a commercially driven, rapidly deployable mission, Twinkle can offer substantial, on-demand telescope time, potentially thousands of hours for each scientific theme. Therefore, stacking multiple observations as availability permits is a viable and efficient strategy for the Twinkle science community, provided that stacking does not introduce significant data distortions.
5 Conclusions
Here we present updates on the Twinkle space missions simulated performance for characterising exoplanetary atmospheres, given the latest design choices of the spacecraft and payload and new insight into the atmospheric composition and structure of select exoplanets provided by JWST’s observations. In addition, we have updated the list of exoplanet candidates suitable for characterisation based on currently confirmed planets and TOIs, and incorporating updated planetary and stellar parameters as available in the literature.
More specifically, leveraging recent discoveries from JWST, we have provided more realistic spectral simulations across various exoplanet types and assessed the retrievability of their atmospheric parameters and chemical abundances at different achievable SNR levels. Retrieval analyses for HD 209458 b, WASP-107 b, and GJ 3470 b have indicated how increased observational investment to reach higher SNR improves the retrievability of the atmospheric parameters and chemical species for different planetary categories. At Twinkle’s native resolution, already at low median SNR values of 3 or 5, key atmospheric properties, such as temperature and cloud altitudes, are well constrained, alongside major molecular species. Increasing the median SNR to 7 or 10 further improves constraints on minor species, exemplified by the successful retrieval of SO2 in GJ 3470 b. For small planets with atmospheres heavier than hydrogen, such as 55 Cnc e, emission retrieval studies revealed that after stacking ten eclipses, temperature structure and dominant atmospheric constituents could be constrained. Additional eclipse observations showed limited improvement, however 55 Cnc e would be an excellent candidate for phase-curve observations with Twinkle. This example showcases that tailored observational strategies are necessary for different exoplanet types to ensure an efficient use of the telescope time.
Our analysis indicates that approximately 100 exoplanets are suitable for in depth atmospheric studies and over 300 exoplanets are good candidates for characterisation.
Future studies will build upon this work to better inform the observational planning of the Twinkle’s exoplanet survey. In particular we plan to include more realistic assumptions of target observability due to Earth obstruction and quantify the impact of stellar activity in our simulations. We also plan to refine our simulations by integrating more advanced cloud formation models and to assess the detectability of other atmospheric features. Additionally, the radiometric model will be continuously updated alongside instrumental or operational developments to ensure precise noise estimates. We will also broaden the scope of simulations to include phase-curve observations, exploring expected outcomes for other scientific themes and their mutual refinements.
In the next five years, numerous space- and ground-based observatories dedicated to exoplanetary science research will become operational. BSSL-Twinkle uniquely complements these efforts by offering the science team full flexibility in the selection of the list of targets and observational strategy. To support informed decision-making by the Twinkle science team, BSSL is committed to clearly showcasing the telescope’s anticipated scientific outcomes.
Acknowledgements
We gratefully acknowledge insightful discussions and inputs from members of the Twinkle Science Team, and their participation in Twinkle Science Conference.
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author.
Appendix A Retrieval Posterior Plots
The posterior plots presented in this section of Appendix illustrate the statistical distributions of retrieved atmospheric parameters obtained from our retrieval analyses. Each posterior plot summarises the likelihood distributions of key atmospheric properties, including planetary radius, equilibrium temperature, cloud-top pressure, and molecular abundances, under various observational conditions (e.g., different SNR levels). The contours indicate parameter uncertainties and correlations. Narrower, well-defined distributions imply robust retrievals and tightly constrained parameters, whereas broader, elongated distributions reflect higher uncertainties or parameter degeneracies. These visualisations highlight how observational strategies, such as varying SNRs or the number of stacked transits, impact the precision and reliability of atmospheric characterisation.
Appendix B Mass Estimates for TESS Object of Interests
In order to estimate masses for TOIs with radius available, we utilised Forecaster tool developed by Chen & Kipping (2017), a code that predict mass from radius for objects. We take the maximum radius of known exoplanets as an edge, and treat those with larger radii as dwarfs and ignore them in this study. Mass estimate using Forecaster tool for TOIs with radius less than 8 Earth radii is being adopted as they align with the known exoplanets distribution as shown in Figure 15. For TOIs with higher radius, we assume an asymmetric log-normal distribution to randomly set the mass, with the mean value set as the median value of known exoplanets with radius larger than 8 Earth radii. Resultant mass distribution for TOIs in Twinkle’s FOR is shown in Figure 15.

### v2

@@
-Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations–B Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations
+Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations–D Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations
 Abstract
@@
 Twinkle is developed, managed and operated by Blue Skies Space (BSSL), a space science data company whose vision is to accelerate and expand the availability of new, high-quality datasets to researchers worldwide, complementing the space-observatories delivered by government space agencies.
-Over its life-time, Twinkle will conduct large-scale survey programs. The scientific objectives and observational strategy of these surveys are defined by researchers who join the Science Team.
+Over its life-time, Twinkle will conduct large-scale survey programs.
+The scientific objectives and observational strategy of these surveys are defined by researchers who join the Science Team.
 Leveraging advances made possible by recent observations with the James Webb Space Telescope, we present here updated simulations evaluating Twinkle’s observational capabilities in the context of exoplanet atmospheres.
@@
 As of 12 March 2025, TESS has identified 7,525 TESS Objects of Interest (TOIs).
-After excluding false positives, 5,253 candidates remain confirmed by the TESS Follow-up Observing Program Working Group (TFOPWG) (Akeson & Christiansen, 2019).
+After excluding false positives, 5,253 candidates remain confirmed by the TESS Follow-up Observing Program Working Group (TFOPWG) (Akeson and Christiansen, 2019).
 This wealth of discoveries of transiting planets around bright stellar hosts has been a game changer for atmospheric characterisation studies using transit, eclipse and phase-curve spectroscopy.
@@
 Ariel aims to conduct a comprehensive chemical survey of exoplanetary atmospheres using transit spectroscopy over the spectral range of 0.5–7.8 m, thereby deepening our understanding of how atmospheric chemistry correlates with planetary characteristics (Tinetti et al., 2018, 2022).
-The Chinese Space Station Telescope (Xuntian), scheduled for late 2026, will offer high-resolution multiband imaging and slitless spectroscopy across 0.255–1 m (Zhan, 2021).
 The Twinkle Space Telescope, managed and operated by Blue Skies Space Ltd.
-(BSSL), will be dedicated to provide spectroscopic data for various science themes developed by its members (Jason et al., 2016; Savini et al., 2018; Edwards et al., 2019a, b; Archer et al., 2020; Stotesbury et al., 2022, 2024).
+(BSSL), will be dedicated to provide spectroscopic data for various science themes developed by its members (Jason et al., 2016; Savini et al., 2018; Edwards et al., 2019c, a; Archer et al., 2020; Stotesbury et al., 2022, 2024).
 Equipped with a 0.45-metre telescope and a spectrometer covering simultaneously from 0.5 to 4.5 m, Twinkle is capable of providing spectra for a wide range of targets such as brown dwarfs, stars, protoplanetary disks, Solar System objects and exoplanets.
@@
 Previous studies by Edwards et al.
-(2019c) evaluated Twinkle’s expected performance to study the atmospheric composition and structure of three iconic exoplanets, HD 209458 b, GJ 3470 b, and 55 Cnc e.
+(2019b) evaluated Twinkle’s expected performance to study the atmospheric composition and structure of three iconic exoplanets, HD 209458 b, GJ 3470 b, and 55 Cnc e.
 However, due to lack of IR spectroscopic data pre-JWST, those analyses were based on theoretical models and simplified assumptions.
@@
 (2024) have identified the presence of several molecular species in the atmosphere of this planet, including H2O, CH4, CO, CO2, SO2, and NH3 when combining HST and JWST data.
-| HD 209458 b | WASP-107 b | GJ 3470 b | 55 Cancri e | |||||||||
-| Planet Type | Hot Jupiter | Warm Neptune | Warm Neptune | Super Earth | ||||||||
-| Planet Mass () | ||||||||||||
-| Planet Radius () | ||||||||||||
-| Teq (K) | ||||||||||||
-| Stellar Magnitude (V/K) |
-| JWST Availability (m) | 2.3–4.5 | 2.4–4.5 | 2.45–4.97 | 3.94–12 | ||||||||
-| HST Wavelength Range (m) | 0.5–0.9; 1.1–1.7 | 0.8–1.6 |
+| HD 209458 b | WASP-107 b | GJ 3470 b | 55 Cancri e | |
+| Planet Type | Hot Jupiter | Warm Neptune | Warm Neptune | Super Earth |
+| Planet Mass () | ||||
+| Planet Radius () | ||||
+| Teq (K) | ||||
+| Stellar Magnitude (V/K) | 7.65 0.03 (V) 6.308 0.026 (K) | 11.592 0.208 (V) 8.637 0.023 (K) | 12.332 0.016 (V) 7.989 0.023 (K) | 5.95 0.023 (V) 4.015 0.036 (K) |
+| JWST Availability (m) | 2.3–4.5 | 2.4–4.5 | 2.45–4.97 | 3.94–12 |
+| HST Wavelength Range (m) | 0.5–0.9; 1.1–1.7 | 0.8–1.6 | 0.524–1.027; 1.1–1.7 | – |
 This paper first reviews Twinkle’s instrumental capability, the models and methodology used in estimating Twinkle’s science performance in Section 2.
@@
 Channel 0 operates from 0.5 to 2.43 m with a resolving power () of up to 70, and channel 1 operates from 2.43 to 4.5 m with a resolving power of up to 50.
-Twinkle’s radiometric performance is assessed using the Twinkle Radiometric Tool (Stotesbury et al., 2022), adapted from the open-source radiometric software ExoRad (Mugnai et al., 2023).
+Twinkle’s radiometric performance is assessed using the Twinkle Radiometric Tool (Stotesbury et al., 2022), adapted from the open-source radiometric software ExoRad2 (Mugnai et al., 2023).
 The radiometric model simulates observational and instrumental noise contributions, including photon noise, instrument emission, detector dark current, read noise, and zodiacal background.
-This model has been refined to incorporate recent advances in instrument design.
-It calculates realistic noise estimates, which in turn allows determination of the number of observations necessary to reach a specific SNR through observation stacking.
+This model has been refined to incorporate recent advances in instrument design (as elaborated in Stotesbury et al.
+(2024)).
+Detailed noise contributions for each target simulated in this study are presented in Appendix A.
+It is worth noting that the scope of this study covers only radiometric simulation, without detailed time-domain simulation, which will be addressed in a forthcoming paper.
+However, an average jitter value derived from ground tests is applied.
+The assumed instrument parameters are summarised in Table 9 of Appendix A.
+The radiometric tool calculates realistic noise estimates, which in turn allows determination of the number of observations necessary to reach a specific SNR through observation stacking.
 The cumulative SNR achieved by combining multiple observations can be expressed as:
-where is the number of observations, and is the signal-to-noise ratio (SNR) of a single observation.
+where is the number of observations (each corresponding to a single transit or eclipse event), and is the signal-to-noise ratio (SNR) of a single observation.
 The signal is derived from the amplitude () of spectral features in transit spectra, approximated by:
@@
 where is Boltzmann’s constant, the equilibrium temperature of the planet, the planet’s surface gravity, and the mean molecular weight of the atmosphere (usually assumed as 2.3 for H2/He atmospheres but refined using recent JWST data in this study).
-Previously, due to limited observational constraints, atmospheric thickness () was set as 1, 3, or 5 scale heights, with 5 scale heights being a commonly used assumption (e.g., Edwards et al., 2019c; Booth et al., 2024).
+Previously, due to limited observational constraints, atmospheric thickness () was set as 1, 3, or 5 scale heights, with 5 scale heights being a commonly used assumption (e.g., Edwards et al., 2019b; Booth et al., 2024).
 However, with the aid of recent data release from JWST and its overlapping wavelength coverage with Twinkle, constraints on atmospheric composition have improved, enabling more precise estimates of atmospheric scale height and thickness above the optically thick continuum.
@@
 However, observational availability of targets are influenced by Earth obstruction due to Twinkle’s orbital configuration, causing Twinkle to sometimes capture partial transits or eclipse events, similar to the situation encountered by CHEOPS (Benz et al., 2021).
-While for some exoplanets, the number of required observations is high, it is important to note that lowering spectral resolution can increase the SNR per wavelength bin and thus reduce the observational requirements.
-Although this method is widely used (e.g., Kreidberg et al., 2014; Stevenson et al., 2014), artificially reducing resolution in simulations introduces bias along the wavelength axis, particularly around spectral peaks.
-To prevent this bias, this study retrieves spectra at Twinkle’s full resolution.
-After the launch of Twinkle, the actual resolution adjustment will effectively improve this bias.
+The extent of transit or eclipse coverage depends on the target’s declination, the satellite’s orbital phase, and the ratio of the planet’s orbital period to Twinkle’s orbital period.
+Partial transits can still yield scientifically useful data but at reduced observing efficiency compared with full-coverage events.
+A detailed quantification of this efficiency factor, incorporating the effects of Earth obstruction and target-specific orbital geometry, will be addressed in a forthcoming time-domain simulation study.
+While for some exoplanets, the number of required observations is high, it is important to note that lowering spectral resolution can increase the SNR per wavelength bin and thus achieve the desired SNR with fewer stacked observations.
+Although post-observation rebinning is widely used (e.g., Kreidberg et al., 2014; Stevenson et al., 2014), artificially reducing resolution in simulations introduces a loss of spectral information along the wavelength axis, particularly around spectral peaks.
+To avoid this, this study retrieves spectra at Twinkle’s full native resolution.
+After launch, rebinning applied to real observational data—rather than to simulated spectra—will mitigate this limitation, as it preserves the full information content of the original observations.
 2.2 Twinkle’s Target Candidates
 Twinkle will be placed in a 1200 km, sun-synchronous low-Earth orbit, with a 6am Local Time of Ascending Node (LTAN).
-The spacecraft’s observational capabilities are restricted by the Field of Regard (FoR) (a 40 degree cone around the anti-Sun vector), solar & lunar exclusion constraints and Earth obstruction across different seasons.(Stotesbury et al., 2022).
+The spacecraft’s observational capabilities are restricted by the Field of Regard (FoR) (a 40 degree cone around the anti-Sun vector), solar & lunar exclusion constraints and Earth obstruction across different seasons (Stotesbury et al., 2022).
 To systematically assess target observability, we employed the BSSL Twinkle Radiometric Tool to simulate observational noise for all confirmed exoplanets within Twinkle’s FoR possessing well-defined stellar and planetary parameters.
 Planetary and stellar data from the NASA Exoplanet Archive are taken as input into the radiometric tool to simulate noise levels for single transits or eclipses at Twinkle’s maximum resolving power.
-For TESS Objects of Interest (TOIs) lacking direct planetary mass measurements and thus not able to be processed with Radiometric Tool, masses were estimated using the methodology described by Chen & Kipping (2017).
+For TESS Objects of Interest (TOIs) lacking direct planetary mass measurements and thus not able to be processed with Radiometric Tool, masses were estimated using the Forecaster tool (Chen and Kipping, 2017).
 Approximately 200 TOIs within Twinkle’s FoR were subsequently excluded due to unavailable stellar mass data.
-Additionally, initial mass-radius relationships for gas giants between 8 and 22 Earth radii proved inconsistent with current empirical data.
-Hence, planetary masses for these TOIs were recalculated based on a log-normal distribution fitted to known transiting exoplanet populations, as detailed in Figure 15 and further described in Appendix B.
+Additionally, planets with radii between 8 and 22 Earth radii fall outside the calibrated range of Forecaster, and the initial mass-radius relationships for gas giants proved inconsistent with current empirical data.
+Hence, planetary masses for these TOIs were recalculated based on a log-normal distribution fitted to known transiting exoplanet populations, as detailed in Figure 16 and further described in Appendix C.
+TOIs lacking sufficient stellar or planetary data for reliable mass estimation were excluded from the candidate list.
 Our primary focus was on exoplanets with H2/He dominated atmospheres; thus, we assumed a mean molecular weight of 2.3 for initial assessments.
@@
 Exoplanets predicted to achieve median SNR levels of 3, 5, or 7 within 30 transits or eclipses at Twinkle’s native spectral resolution were considered suitable candidates for detailed atmospheric characterisation.
+The threshold of 30 observations represents an assumed upper limit per target, determined by Twinkle’s total available observing time, the number of targets in the survey, and the accessibility constraints imposed by the satellite’s orbit over the mission lifetime.
+This choice of upper limit is an operational planning decision rather than a constraint imposed by Twinkle’s instrumental capabilities.
 2.3 Simulated Exoplanets and Spectral Retrievals
@@
 Stellar spectra for each target star were simulated using the PHOENIX stellar library (Husser et al., 2013).
-Chemical profiles were generated using the FastChem equilibrium chemistry plugin for equilibrium assumptions, supplemented by TauREx’s free chemistry feature for non-equilibrium chemistry scenarios.
+Chemical profiles were generated using the FastChem equilibrium chemistry plugin (Stock et al., 2018) for equilibrium assumptions, supplemented by TauREx’s free chemistry retrieval mode.
+In this mode, the volume mixing ratio of each molecule is treated as a free parameter, constant with altitude, and retrieved independently without imposing chemical equilibrium constraints (Al-Refaie et al., 2021).
+While this approach does not explicitly model disequilibrium chemistry, it provides a model-agnostic framework that can capture departures from equilibrium, such as those arising from vertical quenching or photochemistry.
 All atmospheric models assumed a plane-parallel geometry comprising 100 layers, including molecular absorption, Rayleigh scattering, grey clouds, and collision-induced absorption (CIA).
@@
 Retrieval analyses were performed using the Bayesian MultiNest algorithm (Feroz et al., 2009), employing 1000 live points and an evidence tolerance threshold of 0.5.
+Table 2 summarises the prior distributions adopted for all retrieval parameters.
 Non-informative priors for molecular abundances were uniformly set on a log scale within the range of to , while other planetary parameters utilised priors constrained by previous literature of specific planets.
+| Parameter | Prior type | Range |
+| Common to all planets | ||
+| All molecular log-VMRs | LogUniform | |
+| HD 209458 b | ||
+| (K) | Uniform | |
+| factor | Uniform | |
+| (Pa) | LogUniform | |
+| WASP-107 b | ||
+| (K) | Uniform | |
+| factor | Uniform | |
+| (Pa) | LogUniform | |
+| GJ 3470 b | ||
+| (K) | Uniform | |
+| factor | Uniform | |
+| (Pa) | LogUniform | |
+| 55 Cnc e | ||
+| (K) | Uniform | |
+| (K) | Uniform | |
+| factor | Uniform | |
+| (Pa) | LogUniform | |
+| (Pa) | LogUniform | |
+To quantify the detectability of individual molecular species, we employ the Likelihood Ratio Test (LRT) following the methodology described by Tessenyi et al.
+(2013).
+To guard against unreliable MAP estimates for poorly constrained parameters, we perform the LRT using both the MAP and posterior median values; a molecule is classified as a “Detection” only if both tests independently exceed the threshold.
+The full LRT procedure is detailed in Appendix D.
+The more conservative (lower) detection significance is reported in the results tables; molecules below the threshold are classified as “Upper Limit”.
 To incorporate recent JWST observational findings, chemical compositions and planetary parameters for each simulated exoplanet were based on best-fit solutions reported in recent literature, alongside equilibrium chemistry predictions.
@@
 (2025).
-The simulated spectrum is fitted with observed data through minor adjustments.
+The simulated spectrum was aligned with the observed data by applying a vertical offset to reconcile differences in absolute transit depth calibration between the HST and JWST instruments at overlapping wavelengths.
+Molecular abundances and temperature profile parameters were then fine-tuned to optimise the fit to the combined HST+JWST dataset.
 Similarly, to assess Twinkle’s capabilities for WASP-107 b, a simulated spectrum covering Twinkle’s wavelength coverage (0.5–4.5 m) was constructed using HST and JWST NIRSpec data as benchmarks.
@@
 The baseline scenario employed the previously established retrieval setup for WASP-107 b, stacking six transits to achieve a median SNR of 10.
-Individual retrieval analyses were then performed separately for each molecule, progressively increasing their abundances, to quantify their retrieval thresholds under varying observational conditions.
+For each molecule, the abundance was scaled from its baseline value by amplification factors ranging from 5 to 80 (the specific factors tested for each species are listed in Table 8), while all other atmospheric parameters were held fixed at their baseline values.
+Independent retrieval analyses were then performed at both SNR = 5 and SNR = 10 to determine the minimum amplification factor required for a constrained retrieval of each species.
 3 Results
@@
 These preliminary outcomes, integrating JWST observational results, will continue to evolve as additional JWST data become available.
-| Planet | Median SNR | Predicted Transits (5) | Predicted Transits (3) | Predicted Transits (2) | Realistic Transits |
-| HD 209458 b | 5 | 1 (0.18) | 1 (0.51) | 2 (1.21) | 1 (0.85) |
-| 7 | 1 (0.35) | 2 (1.04) | 3 (2.6) | 2 (1.77) | |
-| 10 | 1 (0.75) | 3 (2.32) | 7 (6.79) | 5 (4.24) | |
-| WASP-39 b | 5 | 3 (2.26) | 7 (6.31) | 15 (14.31) | 3 (2.11) |
-| 7 | 5 (4.45) | 13 (12.45) | 29 (28.45) | 5 (4.14) | |
-| 10 | 10 (9.12) | 26 (25.74) | 60 (59.84) | 9 (8.48) | |
-| WASP-80 b | 5 | 3 (2.01) | 6 (5.7) | 14 (13.41) | 8 (7.49) |
-| 7 | 4 (3.98) | 12 (11.56) | 29 (28.5) | 16 (15.35) | |
-| 10 | 9 (8.33) | 26 (25.44) | 71 (70.86) | 35 (34.67) | |
-| WASP-107 b | 5 | 1 (0.16) | 1 (0.45) | 2 (1.02) | 2 (1.31) |
-| 7 | 1 (0.32) | 1 (0.88) | 2 (2.00) | 3 (2.59) | |
-| 10 | 1 (0.65) | 2 (1.81) | 5 (4.15) | 6 (5.4) | |
-| GJ 3470 b | 3 | 1 (0.59) | 2 (1.65) | 4 (3.77) | 4 (3.58) |
-| 5 | 2 (1.64) | 5 (4.69) | 12 (11.09) | 11 (10.49) | |
-| 7 | 4 (3.27) | 10 (9.55) | 24 (23.85) | 23 (22.45) |
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 The realistic observation requirements based on recent JWST constraints are marked in red in Table 3, while atmospheric thickness predictions closest to the JWST-derived reality are highlighted in green.
@@
 Further refinement of the candidate list will be informed by complementary observations from JWST, PLATO, CSST, and Earth 2.0, enabling tailored observational strategies aligned with the scientific objectives of the Twinkle Science Team.
+The candidate list presented in this work will be made publicly available to support the broader exoplanet community in planning complementary observations.
 3.3 Spectral Retrievals: Gas Giants
@@
 (2024) was used for these molecules in a free retrieval analysis.
-Results in Figure 11 confirm robust retrieval of H2O, TiO, and VO, whereas other molecular signatures remain masked by the dominant H2O absorption and cloud opacity.
-Twinkle’s extensive spectral coverage from optical to infrared allows for clear identification of TiO and VO absorption features.
+Results in Figure 12 confirm robust retrieval of H2O and TiO, whereas other molecular signatures remain masked by the dominant H2O absorption and cloud opacity.
+Twinkle’s extensive spectral coverage from optical to infrared allows for clear identification of TiO absorption features, while VO remains below the detection threshold (LRT ; Table 4) despite showing a posterior shift from the prior.
 However, due to the relatively low assumed abundance of CO2 () and increased noise at longer wavelengths, the CO2 feature was only partially detected and thus less reliably constrained.
@@
 This limitation of improving SNR and difficulty of retrieving minor species in HD 209458 b likely arises from spectral masking by prominent water absorption and simplified grey cloud assumptions, rather than reflecting Twinkle’s intrinsic capabilities.
-| Parameter | Truth | Retrieved (SNR = 5) |
-| Planet Radius [] | 1.35 | |
-| Temperature [K] | 1140 | |
-| Detection | ||
-| Upper Limit | ||
-| 2.11 | ||
+| Parameter | Truth | Retrieved (SNR = 5) | LRT () |
+| Planet Radius [] | 1.35 | ||
+| Temperature [K] | 1140 | ||
+| Detection | |||
+| 15.18 | |||
+| 8.99 | |||
+| Upper Limit | |||
+| 2.05 | |||
+| 1.20 | |||
+| 0.10 | |||
+| 0.05 | |||
+| 0.12 | |||
+| 2.11 | |||
 3.3.2 WASP-107 b
@@
 Furthermore, while CO2 could be constrained at all simulated SNR levels, increasing the SNR to 7 or 10 significantly improved the accuracy and reliability of its retrieval.
-| Parameter | Truth | RP (SNR = 5) |
-| Planet Radius [] | 0.933 | |
-| [K] | 700 | |
-| Detection | ||
-| Upper Limit | ||
+| Parameter | Truth | RP (SNR = 5) | LRT () |
+| Planet Radius [] | 0.933 | ||
+| Temperature [K] | 700 | ||
+| Detection | |||
+| 13.39 | |||
+| Upper Limit | |||
+| 1.35 | |||
+| 0.07 | |||
+| 0.08 | |||
+| 0.01 | |||
 3.3.3 GJ 3470 b
@@
 Conversely, molecules like SO2, with a single prominent absorption feature near the spectral range boundary, require high SNR levels for accurate abundance retrieval.
-| Parameter | Truth | RP (SNR = 3) |
-| Planet Radius [] | 0.36 | |
-| Temperature [K] | 600 | |
-| Detection | ||
-| Upper Limit | ||
+| Parameter | Truth | RP (SNR = 3) | LRT () |
+| Planet Radius [] | 0.36 | ||
+| Temperature [K] | 600 | ||
+| Detection | |||
+| 5.03 | |||
+| 3.37 | |||
+| 9.52 | |||
+| Upper Limit | |||
+| 0.02 | |||
+| 0.39 | |||
 3.4 Spectral Retrievals: super-Earths
@@
 Twinkle’s expected performance, calculated by reducing the resolution of channel 0 to one-third and channel 1 to half their native resolutions and stacking ten eclipse observations, demonstrates comparable sensitivity to JWST within its spectral range.
-A combined analysis of Figures 7 and 8 reveals significant spectral divergence from blackbody emission beginning around 2.0 m.
+Figure 7 reveals significant spectral divergence from blackbody emission beginning around 2.0 m.
 Although the higher resolution of channel 0 below this wavelength is underutilised, the spectral range above 2.0 m effectively captures prominent CO2 absorption features.
 Additional CO2 features are clearly identifiable within Twinkle’s channel 1 coverage.
-Figure 8 further highlights that CO2 dominates the spectral signature throughout Twinkle’s wavelength range, overshadowing contributions from minor species such as C2H2, CO, and HCN.
+The contribution curves further highlight that CO2 dominates the spectral signature throughout Twinkle’s wavelength range, overshadowing contributions from minor species such as C2H2, CO, and HCN.
 Notably, an increase in CO contribution is evident beyond 4.2 m, suggesting enhanced detectability if Twinkle’s coverage could extend slightly further.
 Using this simulated spectrum as input, we performed emission retrievals to evaluate Twinkle’s capability in constraining atmospheric properties.
-To prevent spectral bias from wavelength binning, retrievals were conducted at Twinkle’s native resolution, despite larger resultant uncertainties.
+To avoid the loss of spectral information introduced by post-processing wavelength binning, retrievals were conducted at Twinkle’s native resolution, despite larger resultant uncertainties.
 Retrieval outcomes for scenarios involving 10, 20, and 30 stacked eclipse observations are presented in Figure 15.
@@
 Additional observations enhance parameter constraints incrementally but do not significantly improve the detection of dominant atmospheric constituents.
-This demonstrates Twinkle’s robust performance, particularly for bright terrestrial targets.
-| Parameter | Truth | RP (10 eclipses) |
-| Planet Radius [] | 0.167 | |
-| [K] | 2300 | |
-| [K] | 1500 | |
-| [Pa]) | 5 | |
-| [Pa]) | 2 | |
-| Detection | ||
-| Upper Limit | ||
+This demonstrates Twinkle’s robust performance, particularly for bright super-Earth targets such as 55 Cnc e.
+| Parameter | Truth | RP (10 eclipses) | LRT () |
+| Planet Radius [] | 0.167 | ||
+| [K] | 2300 | ||
+| [K] | 1500 | ||
+| [Pa]) | 5 | ||
+| [Pa]) | 2 | ||
+| Detection | |||
+| 5.38 | |||
+| Upper Limit | |||
+| 0.01 | |||
+| 0.01 | |||
 3.5 Minor Species Sensitivity Study on WASP-107 b
-Contribution plots from simulations (Figures 5, 6, and 8) revealed that several minor species, such as CH4, SO2, NH3, and H2S, typically possess abundances too low to yield distinguishable spectral features, especially in atmospheres dominated by water vapour absorption or obscured by clouds.
+Contribution plots from simulations (Figures 5, 6, and 7) revealed that several minor species, such as CH4, SO2, NH3, and H2S, typically possess abundances too low to yield distinguishable spectral features, especially in atmospheres dominated by water vapour absorption or obscured by clouds.
 Direct retrieval of these species from featureless continuum regions was thus found unreliable, as illustrated in posterior plots provided in the Appendix.
 Such limitations reflect realistic observational challenges faced when attempting to detect minor constituents even at higher SNR levels.
-To better quantify these observational constraints, we assessed the amplification factors necessary for each minor species’ spectral features to surpass the noise threshold of Twinkle observations at SNR = 10 (Comparison of the original and amplified sepctra in Figure 9).
+To better quantify these observational constraints, we assessed the amplification factors necessary for each minor species’ spectral features to surpass the noise threshold of Twinkle observations at SNR = 10 (Comparison of the original and amplified spectra in Figure 8).
 Our results show that CH4 and SO2 features become clearly observable when their abundances increase by approximately fivefold, corresponding to abundances of and , respectively.
@@
 H2S retrieval was challenging at SNR = 5 due to its weaker, indistinct absorption features compared to water.
+While the amplified spectra may appear visually consistent with the baseline within individual spectral bin error bars (Figure 8), the LRT analysis captures the cumulative effect of systematic spectral differences across multiple bins simultaneously.
+The retrieval posteriors in Figure 9 confirm that the amplified species are indeed constrained, validating the sensitivity thresholds identified by our analysis.
+| Molecule | Amplification | SNR 10 | SNR 5 |
+| SO2 | baseline | ||
+| 5 | 4.69 | 0.12 | |
+| 9 | – | 6.20 | |
+| NH3 | baseline | 0.01 | – |
+| 5 | 0.03 | – | |
+| 10 | 8.23 | – | |
+| 20 | 13.58 | – | |
+| 40 | – | 11.99 | |
+| 50 | 24.10 | – | |
+| H2S | baseline | – | |
+| 5 | 0.17 | – | |
+| 10 | 2.42 | – | |
+| 20 | 3.87 | – | |
+| 50 | 24.44 | – | |
+| 80 | – |
+Table 8 reports the LRT detection significance for each amplification scenario.
+SO2 achieves robust detection () at 5 amplification with SNR 10 and at 9 amplification () with SNR 5.
+NH3 requires 10 amplification at SNR 10, while H2S requires 20.
+CH4 is not included in the table because it remains undetected by the relative LRT (full model versus model without CH4) at all tested amplification levels.
+This non-detection arises from strong spectral degeneracy with H2O, whose retrieved abundance (log VMR ) produces broad near-infrared absorption that overlaps and masks the CH4 features; the retriever can compensate for the removal of CH4 by adjusting H2O, yielding a low LRT statistic even when the CH4 posterior itself is constrained (Figure 9).
+However, an alternative LRT formulation using a featureless (flat-line) null hypothesis confirms that the CH4 spectral signature is intrinsically detectable by BSSL, exceeding at 5 amplification, consistent with the spectral amplification results shown in Figure 8.
+SiO, whose spectral contribution is negligible in all scenarios, is likewise omitted from the table.
 The goal of this sensitivity analysis was to isolate the impact of instrumental capabilities (resolution, SNR) from atmospheric factors such as inherently low abundances and optically thick cloud cover.
@@
 Cloud contributions may significantly impact parameters’ retrieval, especially for minor species.
-For instance, in WASP-107 b, SO2 spectral features were fully obscured by cloud opacity, and there are degeneracies between cloud height and planetary radius.
+For instance, in our simulated retrievals of WASP-107 b, SO2 spectral features were fully obscured by the assumed cloud opacity (see Figure 5), and degeneracies between cloud-top height and planetary radius were observed.
 Incorporating detailed cloud microphysical and radiative models into future simulations might allow more realistic analyses.
@@
 Investigating stellar activity effects on exoplanet atmospheres is one of Twinkle’s core scientific themes, and a future study will explore Twinkle’s capability to identify and quantify these stellar variations.
+In addition to astrophysical variability, instrumental systematics such as detector drift, thermal variations, and pointing jitter may affect the consistency of stacked observations.
+A detailed assessment of these instrumental effects will be presented in a companion study focusing on time-domain simulations.
 Another significant constraint is observational availability, influenced by Earth obstruction due to Twinkle’s orbital configuration, planetary orbital periods, and transit durations.
@@
 This example showcases that tailored observational strategies are necessary for different exoplanet types to ensure an efficient use of the telescope time.
-Our analysis indicates that approximately 100 exoplanets are suitable for in depth atmospheric studies and over 300 exoplanets are good candidates for characterisation.
+Our analysis indicates that approximately 100 exoplanets are suitable for in depth atmospheric studies and over 300 exoplanets are good candidates for characterisation with Twinkle.
 Future studies will build upon this work to better inform the observational planning of the Twinkle’s exoplanet survey.
@@
 We gratefully acknowledge insightful discussions and inputs from members of the Twinkle Science Team, and their participation in Twinkle Science Conference.
+This work made use of data and results from observations with the NASA/ESA/CSA James Webb Space Telescope, as reported in the published literature.
+The JWST data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST.
 Data Availability
 The data underlying this article will be shared on reasonable request to the corresponding author.
-Appendix A Retrieval Posterior Plots
+Conflict of Interest
+The authors declare no conflict of interest.
+Appendix A Noise Contributions
+The noise component plots presented in this section illustrate how the different noise sources considered by the Radiometric Tool vary across each target.
+For all targets simulated in this work, the noise budget remains in the photon-noise-dominated regime, with all noise contributions increasing towards the edges of each spectral channel.
+For relatively bright targets such as HD 209458 b and 55 Cnc e, read noise begins to exceed photon noise at the longer wavelengths covered by channel 1, as the detector requires more reads to avoid saturation.
+The radiometric simulations presented in this study adopt the instrument parameters summarised in Table 9.
+In its current design, the telescope has an effective collecting area of m2.
+The optical train consists of four mirror surfaces and a dichroic beam splitter in the common path, followed by a dispersing prism and three fold mirrors in each channel.
+All mirror surfaces are assumed to have a reflectivity of 98%, while the dichroic and prism each transmit 90% and 80%, respectively.
+The detector quantum efficiency is wavelength-dependent, ranging from 75% to 95% across the 0.5–4.5 m bandpass.
+The point spread function (PSF) in channel 0 is treated as diffraction-limited, while channel 1 assumes a wavefront error of m.
+Pixel-based PSF models are used for both channels.
+The detectors operate at 90 K with a read noise of 25 counts per read, a dark current of 35 counts s-1, and a well depth of 72,000 counts (80% of which is usable to minimise non-linear detector effects).
+For bright targets such as HD 209458 b and 55 Cnc e, shorter integration times are required to avoid well saturation, leading to a higher number of detector reads per exposure and consequently elevated read noise contributions (Figures 10(a) and 11(b)).
+For fainter targets (WASP-107 b, GJ 3470 b), longer integrations are feasible, and photon noise dominates across the full spectral range.
+Line-of-sight pointing jitter is included in the noise model using a pre-computed jitter profile derived from ground-test pointing performance data.
+The out-of-transit noise overhead is accounted for by a factor , corresponding to an out-of-transit baseline equal to 1.5 times the transit duration.
+The zodiacal background is modelled using a position-dependent sky map with a nominal scaling factor of 2.5.
+Additional systematic noise contributions (e.g.
+gain noise, detector drift, thermal variations) are not included in the present radiometric model but will be addressed in a forthcoming time-domain simulation study.
+| Parameter | Value |
+| Collecting area () | 0.146 m2 |
+| Channel 0 range | 0.5–2.43 m () |
+| Channel 1 range | 2.43–4.5 m () |
+| Mirror reflectivity (per surface) | 98% |
+| Dichroic transmission | 90% |
+| Prism transmission | 80% |
+| Detector QE | 75–95% |
+| Read noise | 25 counts |
+| Dark current | 35 counts s-1 |
+| Well depth | 72,000 counts |
+| Detector temperature | 90 K |
+| Optics temperature | 100–160 K |
+| PSF (Ch 0) | Diffraction-limited |
+| PSF (Ch 1) | m |
+The observation stacking equation (Equation 1) assumes that the noise contributions from successive transits or eclipses are uncorrelated.
+While this assumption is appropriate for photon noise and random detector noise, systematic effects such as instrumental drift, thermal cycling, or stellar variability may introduce correlated noise components that do not average down as .
+Experience with HST/WFC3 transit spectroscopy suggests that systematic noise floors of 20–30 ppm can be reached for bright targets after careful detrending (Deming et al., 2013).
+A detailed assessment of Twinkle’s systematic noise floor, incorporating time-domain instrumental effects, will be presented in a companion study.
+Appendix B Retrieval Posterior Plots
 The posterior plots presented in this section of Appendix illustrate the statistical distributions of retrieved atmospheric parameters obtained from our retrieval analyses.
@@
 These visualisations highlight how observational strategies, such as varying SNRs or the number of stacked transits, impact the precision and reliability of atmospheric characterisation.
-Appendix B Mass Estimates for TESS Object of Interests
-In order to estimate masses for TOIs with radius available, we utilised Forecaster tool developed by Chen & Kipping (2017), a code that predict mass from radius for objects.
+Appendix C Mass Estimates for TESS Object of Interests
+In order to estimate masses for TOIs with radius available, we utilised Forecaster tool developed by Chen and Kipping (2017), a code that predict mass from radius for objects.
 We take the maximum radius of known exoplanets as an edge, and treat those with larger radii as dwarfs and ignore them in this study.
@@
 Resultant mass distribution for TOIs in Twinkle’s FoR is shown in Figure 16.
+Appendix D Likelihood Ratio Test for Molecular Detectability
+To quantify the statistical significance of molecular detections, we employ the Likelihood Ratio Test (LRT) following the methodology described by Tessenyi et al.
+(2013).
+For each molecule under consideration, the alternative hypothesis corresponds to the full forward model containing all retrieved species, while the null hypothesis corresponds to the model with the test molecule removed.
+For each hypothesis, Monte Carlo simulations with 100,000 iterations generate synthetic observations by adding Gaussian noise (drawn from the per-bin radiometric uncertainties) to the corresponding model spectrum.
+For each realisation, the likelihood ratio statistic is computed, yielding distributions of (data generated under ) and (data generated under ).
+A molecule is classified as a “Detection” when median() mean() , where is the standard deviation of the null distribution.
+The detection significance, expressed in units of , is reported in the results tables.
+Molecules falling below the threshold are classified as “Upper Limit”, indicating that their abundances are either prior-dominated or only weakly constrained by the data.
+For emission spectra (55 Cnc e), we adopt the same molecule-removal framework: is the full emission model containing all retrieved species, while is the full model with the test molecule removed.
+This approach is more conservative than the original Tessenyi et al.
+(2013) emission formulation (which uses a featureless blackbody as ), because it tests whether a molecule’s spectral signature is distinguishable in the presence of all other atmospheric constituents, rather than merely deviating from a smooth continuum.
+To ensure robustness against unreliable MAP estimates for poorly constrained parameters, we perform the LRT using both the MAP and posterior median values for each molecule.
+A molecule is classified as a “Detection” only if both tests independently exceed the threshold.
+The more conservative (lower) of the two significance values is reported in the results tables.
+Figures 17–20 show the resulting (black) and (green) distributions for each molecule tested.
+The red dashed line indicates the detection threshold, while the blue solid line marks the median of .
+A molecule is detected when the blue line exceeds the red line.